If a 26 ft long ladder rest against a wall. the distance from the base of the wall to the bottom of the ladder is x. The height at which the ladder rests against the wall is x+14. find the distance, x, from the base of the wall to the bottom of the ladder.
Answer by palanisamy(496)

The ladder forms a right triangle with the wall'
So, x^2+(x+14)^2 = 26^2
x^2+x^2+28x+196 = 676
2x^2+28x-480=0
x^2+14x-240=0
(x+24)(x-10)=0
x=-24 or x=10
x cannot be negative
So, x=10